\begin{thebibliography}{STW03}

\bibitem[Cal20]{Calegari}
Danny Calegari.
\newblock 3-manifolds.
\newblock unpublished, 2020.

\bibitem[CLN06]{Chow-Ni-Lu}
Bennett Chow, Peng Lu, and Lei Ni.
\newblock {\em Hamilton's {R}icci flow}, volume~77 of {\em Graduate Studies in
  Mathematics}.
\newblock American Mathematical Society, Providence, RI; Science Press Beijing,
  New York, 2006.

\bibitem[CM19]{Cortiss-Mur}
Jean~C. Cortissoz and Alexander Murcia.
\newblock The {R}icci flow on surfaces with boundary.
\newblock {\em Comm. Anal. Geom.}, 27(2):377--420, 2019.

\bibitem[CR23]{Cortiss-Reyes}
Jean~C. Cortissoz and C\'{e}sar Reyes.
\newblock Classical solutions to the one-dimensional logarithmic diffusion
  equation with nonlinear {R}obin boundary conditions.
\newblock {\em Math. Nachr.}, 296(9):4086--4107, 2023.

\bibitem[Eck07]{Ecker}
Klaus Ecker.
\newblock A formula relating entropy monotonicity to {H}arnack inequalities.
\newblock {\em Comm. Anal. Geom.}, 15(5):1025--1061, 2007.

\bibitem[Gia16]{Gianniotis}
Panagiotis Gianniotis.
\newblock Boundary estimates for the {R}icci flow.
\newblock {\em Calc. Var. Partial Differential Equations}, 55(1):Art. 9, 21,
  2016.

\bibitem[Ham95]{Hamilton}
Richard~S. Hamilton.
\newblock The formation of singularities in the {R}icci flow.
\newblock In {\em Surveys in differential geometry, {V}ol. {II} ({C}ambridge,
  {MA}, 1993)}, pages 7--136. Int. Press, Cambridge, MA, 1995.

\bibitem[Lot12]{Lott}
John Lott.
\newblock Mean curvature flow in a {R}icci flow background.
\newblock {\em Comm. Math. Phys.}, 313(2):517--533, 2012.

\bibitem[Per02]{perelman2002}
Grisha Perelman.
\newblock The entropy formula for the {R}icci flow and its geometric
  applications, 2002.

\bibitem[STW03]{Sesum-Tian}
Natasa Sesum, Gang Tian, and Xiaodong Wang.
\newblock Notes on {P}erelman's paper on the entropy formula for the {R}icci
  flow and its geometric applications.
\newblock unplublished, 2003.

\end{thebibliography}

\documentclass{amsart}
\usepackage[utf8]{inputenc}

\usepackage{amssymb, amsmath, amsbsy} 
\allowdisplaybreaks

\newtheorem{theorem}{Theorem}[section]
\newtheorem{definition}[theorem]{Definition}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{remark}[theorem]{Remark}

\title[Singular Ricci Flows]{Singular Ricci Flows on surfaces with boundary and positive scalar curvature}   
\author{Jean Carlos Cortissoz - Juan José Villamarín} 
\date{October  2023} 

\AtEndDocument{\bigskip{\footnotesize%
  \textsc{Department of Mathematics, Universidad de los Andes, Carrera 1 No. 18A - 10, 111711 Bogotá, Colombia} \par  
  \textit{E-mail address}, \texttt{jcortiss@uniandes.edu.co} \par
  \textit{E-mail address}, \texttt{jj.villamarin@uniandes.edu.co} 
  }}

\begin{document}

\maketitle

\begin{abstract}
We study the subsequential convergence of singular solutions to the Ricci flow with prescribed constant in space geodesic curvature on compact surfaces with boundary. Furthermore, we show that in the particular case of rotational symmetry, this convergence does not depend on the sign of the geodesic curvature of the boundary.
\end{abstract}

\tableofcontents

\section{Introduction}

We will consider the following PDE system on a compact surface with boundary $(M,g)$,
\begin{equation}\label{eq:Ricci flow}
    \begin{cases}
        \frac{\partial g}{\partial t} = -2Rc = -Rg, & \text{in} \hspace{5pt} M\times (0,T),\\
        H( t) = \psi (t) & \text{on} \hspace{5pt} \partial M \times (0,T),\\
        g(0) = g_0& \text{in} \hspace{5pt} M.
    \end{cases}
\end{equation}
Here $R$ denotes the scalar curvature of the metric $g$ (sometimes we will also employ the notation
$R_{g\left(t\right)}$), and $H$ is the geodesic curvature of the boundary.
From now on, we will assume that this is a singular solution to the Ricci flow, meaning that 
the maximal interval of existence
for the solution to this flow $\left(0,T\right)$ satisfies that
$T< \infty$. We shall also assume that $H(t) = \psi(t)$ is a real-valued function bounded in compact time intervals and constant in space, 
that is, it is only time-dependent. 
It is well known then, that in this case, the scalar curvature must blow up, that is 
\[
\limsup_{t\rightarrow T}\left(\sup_{p\in M} \left|R_{g\left(t\right)}\left(p\right)\right|\right)=\infty.
\]
There are some cases in which it is known that the solution blows up at a finite time. For instance, if $\partial R/ \partial N$ remains 
positive as long as the solution exists, or if $H \leq 0$, and $\mathcal{X}(M) > 0$, then the solution is singular \cite{Cortiss-Mur}.\\


We now state the first result obtained in this work.

\begin{theorem}\label{th:1.1}
    Let $M$ be a compact surface with boundary endowed with a Riemannian metric $g_0$. Then, if $g(t)$ is the solution to the Ricci flow 
    (\ref{eq:Ricci flow}) with boundary condition $\psi\left(t\right)\geq 0$, and satisfies that 
    $R_{g\left(t\right)}>0$, then for any sequence $t_i \rightarrow T$ for which the scalar curvature blows up at maximal rate, it holds that
        \begin{equation}\label{eq:uniformized curvature}
        \lim_{i \rightarrow \infty}\frac{R_\text{max}(t_i)}{R_\text{min}(t_i)} = 1,
        \end{equation}    
        where $R_{\text{max}}\left(t\right)=\max_{p\in M}R_{g\left(t\right)}\left(p\right)$ and
        $R_{\text{min}}\left(t\right)=\min_{p\in M}R_{g\left(t\right)}\left(p\right)$.
        
    Furthermore, the corresponding solution $\Tilde{g}(\Tilde{t})$ to the normalized Ricci flow exists for all time, and for the 
    corresponding sequence $\Tilde{t}_n \rightarrow \infty$ the metrics $\Tilde{g}(\Tilde{t}_{n})$ converge smoothly, up to a subsequence, to 
    a metric of constant curvature and totally geodesic boundary.
\end{theorem}

Observe that this is an improvement of Theorem 1.1 in \cite{Cortiss-Mur}. Indeed, here we drop the assumption that the sign of the time derivative of the geodesic curvature is negative. The proof
of this result follows along the same lines of the analogue result in
\cite{Cortiss-Mur}, and is obtained by considering a modification of a Perelman-type entropy functional 
introduced by Lott in \cite{Lott}; see Section \ref{Sec:monotonicity}.\\

Furthermore, it is also possible to prove the following theorem: under the additional hypothesis of rotational symmetry, the subsequential convergence of the Ricci flow does not depend on the sign of the prescribed geodesic curvature $\psi$ of the boundary.

\begin{theorem}\label{th:1.2}
    Let $(M,g_0)$ be the disk with a rotationally symmetric metric $g_0$. Let $g(t)$ be the solution to the Ricci flow (\ref{eq:Ricci flow}) with boundary condition $\psi\left(t\right)$, and assume that $R_{g\left(t\right)}>0$ holds. Then for any sequence $t_i \rightarrow T$ for which the scalar curvature blows up at maximal rate, we have that 
    \begin{equation}\label{uniformized curvature 2}
        \lim_{i\rightarrow \infty}\frac{R_\text{max}(t_i)}{R_\text{min}(t_i)} = 1.
    \end{equation}
    Meaning that the scalar curvature uniformizes along any such sequence of times. Also,
    if $\psi\left(t\right)\leq 0$, regardless of the sign of the scalar curvature, then (\ref{uniformized curvature 2}) holds.
    Moreover, in both cases, the corresponding solution $\Tilde{g}(\Tilde{t})$ to the normalized Ricci flow exists for all time, and for the corresponding sequence $\Tilde{t}_n \rightarrow \infty$ there is a subsequence $\Tilde{t}_{n_k}$ such that $\Tilde{g}(\Tilde{t}_{n_k})$ converges smoothly to a (rotationally symmetric) metric of constant curvature and totally geodesic boundary.
\end{theorem}

Recall that in the case of rotational symmetry and negative prescribed geodesic curvature, it was only known that the normalized Ricci flow exists for all time; see Theorem 1.2 in \cite{Cortiss-Mur}. Hence, Theorem (\ref{th:1.2}) above not only shows that, up to a subsequence, the normalized Ricci flow also converges to a metric of constant curvature and totally geodesic boundary, but that this convergence does not depend on the sign of the geodesic curvature.

\medskip
Our present work leaves open some interesting questions. For instance, if the flow exists for all time, and $\psi\left(t\right)\geq 0$ the
boundary condition becomes unbounded as $t\rightarrow \infty$, could it happen that $\sup_{p\in M}\left|R_{g\left(t\right)}\left(p\right)\right|\rightarrow 0$
as $t\rightarrow \infty$? How does a solution to the Ricci flow behave when $R_{g\left(0\right)}<0$ and
the prescribed geodesic curvature satisfies $\psi\left(t\right)>0$ for all $t>0$?

\medskip
The study of the Ricci flow on surfaces with boundary is not only interesting from a geometric point of view, as
it also has consequences on the behavior of solutions to the Logarithmic Diffusion Equation with nonlinear boundary conditions
(see for instance \cite{Cortiss-Reyes}),
thus the interest in our studies not only lies in possible geometric applications, but also in their
possible analytic consequences.

\medskip
This paper is organized as follows. In Section \ref{Sec:monotonicity} we introduce the modified Perelman-type
functional and show its monotonicity 
under (\ref{eq:Ricci flow}) given some
assumptions; in Section \ref{Sec:proof1.1} we give a proof of Theorem (\ref{th:1.1}); in Section \ref{Sec:proof1.2}
we give a proof of Theorem (\ref{th:1.2}).


\section{Monotonicity of the Modified Perelman´s Entropy Functional}\label{Sec:monotonicity}

This section is devoted to show that under the Ricci flow and the nonnegativity of the (prescribed) geodesic curvature of the boundary we have a monotonicity formula for the modified Perelman's entropy functional, which is defined below.

\medskip
Before we proceed, 
let us introduce the notation used throughout the document. For a given Riemannian metric $g$ on $M$, let $\nabla$ denote the
corresponding Levi-Civita connection and $\hat{\nabla}$ the induced connection on the boundary $\partial M$. The scalar second fundamental 
form of the boundary will be denoted by $h$. It will be computed with respect to the inward pointing unit normal $\eta$, as well as the 
geodesic curvature $H$. On the other hand, the outward pointing unit normal will be denoted by $N$. Recall that under the Ricci flow, the 
geodesic curvature satisfies the following evolution equation 
\begin{equation*}
    \frac{\partial R}{\partial N} = H R - 2H',
\end{equation*}
see \cite{Cortiss-Mur} for a proof of this. Additionally, $d\mu$ will denote the volume density of $M$, while $d\sigma$ will denote that of $\partial M$.\\

\begin{definition}
    Let $\mathcal{W}_\infty$ be the functional defined by 
    \begin{equation}
        \mathcal{W}_\infty:= \frac{1}{4\pi\tau} \int_M [\tau(R + |\nabla f|^2) + f-2]e^{-f}d\mu + \frac{1}{4\pi}\int_{\partial M}2H e^{-f}d\sigma.
    \end{equation}
\end{definition}

\begin{remark}
    \begin{enumerate}
        \item Observe this is a modification of Lott's $I_\infty$ functional defined in \cite{Lott}. In fact, we have
    \begin{equation*}
        \mathcal{W}_\infty = \frac{1}{4\pi}I_\infty + \frac{1}{4\pi \tau} \int_M (f-2)e^{-f}d\mu.
    \end{equation*}
    \item Furthermore, a similar type of functional has also been used by Ecker in \cite{Ecker}. Indeed, he considered the functional 
    \begin{equation*}
        \mathcal{W}_\beta = \int_\Omega (\tau| \nabla f|^2 + f - n) e^{-f}(4\pi\tau)^{-n/2}d\mu + 2\tau \int_{\partial \Omega}\beta e^{-f/2}(4\pi\tau)^{-n/2}d\sigma,
    \end{equation*}
    where $f$ and $\beta$ are smooth functions on $\Omega$ and $\partial \Omega$ respectively. Note that when $\beta = H$, the mean curvature of the boundary, then $\mathcal{W}_H$ is the functional $\mathcal{W}_\infty$ defined above in the case $R = 0$.
    \end{enumerate}
\end{remark}
Let us now compute the variation of this functional under the following assumptions. To do this, we follow \cite{Sesum-Tian}. First, let us denote 
\begin{equation*}
\frac{\partial}{\partial s} g_{ab} = v_{ab}, \text{ and }\hspace{5pt} \frac{\partial}{\partial s} f = \Tilde{f},
\end{equation*}
and suppose the measure $dm: = e^{-f}d\mu$ is fixed \textit{i.e.} $\frac{\partial}{\partial s}dm = 0$. Therefore, 
\begin{equation*}
    \Tilde{f} = \frac{1}{2}V := \frac{1}{2} v^{a}{}_a.
\end{equation*}
Thus, we simply need to compute the variation of $R + |\nabla f|^2$.
\begin{align*}
    \frac{\partial}{\partial s}(R + |\nabla f|^2) &= -\Delta V + \nabla^{a}V\nabla_a f + \nabla^{a}\nabla^bv_{ab} - v^{ab}(R_{ab} + \nabla_af \nabla_bf) .  
\end{align*}
Multiplying by $e^{-f}$ we obtain the following equality:
\begin{align*}
    e^{-f}(R + |\nabla f|^2) & = -v^{ab}(R_{ab} + \nabla_af\nabla_bf)e^{-f} \\
    &\phantom{=\ }- \nabla_a((\nabla^{a}V)e^{-f}) + (\nabla^{a}\nabla^bv_{ab})e^{-f}\\
    &= -v^{ab}(R_{ab} + \nabla_af\nabla_bf)e^{-f}- \nabla_a((\nabla^{a}V)e^{-f})\\
    & \phantom{=\ } + \nabla^{a}((\nabla^bv_{ab})e^{-f}) - \nabla^bv_{ab}\nabla^{a}e^{-f}\\
    &=  -v^{ab}(R_{ab} + \nabla_af\nabla_bf)e^{-f}- \nabla_a((\nabla^{a}V)e^{-f})\\
    & \phantom{=\ } + \nabla^{a}((\nabla^bv_{ab})e^{-f}) - \nabla^b(v_{ab}\nabla^{a}e^{-f}) + v_{ab}\nabla^{a}\nabla^be^{-f}\\
    &= -v^{ab}(R_{ab} + \nabla_a\nabla_bf)e^{-f}- \nabla_a((\nabla^{a}V)e^{-f})\\
    & \phantom{=\ } + \nabla^{a}((\nabla^bv_{ab})e^{-f} - v_{ab}\nabla^be^{-f}).
\end{align*}
Combining the calculations above and the contributions of the $f-2$ and boundary terms, we obtain the variation of the functional.
\begin{proposition}
    Under the hypotheses above, the variation of the functional reads as follows.
    \begin{equation}\label{eq:variation}
        \begin{split}
            \frac{\partial}{\partial s}\mathcal{W}_\infty &= \frac{1}{4\pi}\int_{M}-v^{ab}(R_{ab} + \nabla_a\nabla_bf - \frac{1}{2\tau}g_{ab})e^{-f}d\mu\\
            & \phantom{=\ }+ \frac{1}{4\pi}\int_{\partial M}\left(2\frac{\partial H}{\partial s} - H v^2{}_2\right)e^{-f}d\sigma - \frac{1}{4\pi}\int_{\partial M}\frac{\partial V}{\partial N}e^{-f}d\sigma\\
            & \phantom{=\ } + \frac{1}{4\pi}\int_{\partial M}(\nabla_bv^{ab} + v^{ab}\nabla_bf)N_ae^{-f}d\sigma,
        \end{split}
    \end{equation}
    where $N$ represents the outward pointing unit normal on $\partial M$.
\end{proposition}

Now, let us define $v_{ab} = -2(R_{ab} + \nabla_a\nabla_bf - \frac{1}{2\tau}g_{ab})$. Hence, we have that 
\begin{equation*}
    V = -2 R - 2\Delta f + \frac{2}{\tau}.
\end{equation*}
In order to prove monotonicity, we proceed by computing the integrand of the last term in Equation (\ref{eq:variation}). To this end, we perform the calculations in a coordinate system $(x^1, x^2)$ such that $x^2 = 0$ is a local defining function for $\partial M$, $\{ \partial_1, \partial_2\}$ is an orthogonal frame and at some fixed (and arbitrary) time $t$ we have $\partial_2 = N$ on $\partial M$. Therefore,
\begin{align*}
    \nabla_b v^b{}_2 + v^b{}_2\nabla_b f &= \nabla_b (-Rg^b{}_2 - 2\nabla^b\nabla_2f + \frac{1}{\tau}g^b{}_2) + v^1{}_2 \nabla_1f + v^2{}_2\nabla_2f\\
    &= -\nabla_2R - 2\Delta \nabla_2 f \\
    &\phantom{=\ }+ (-Rg^1{}_2 - 2 \nabla^1\nabla_2f + \frac{1}{\tau}g^1{}_2)\nabla_1f + v^2{}_2\nabla_2f\\
    &= - \nabla_2R - 2\nabla_2\Delta f - 2 R^b{}_2\nabla_bf\\
    &\phantom{=\ }-2\nabla^1\nabla_2f\nabla_1f + v^2{}_2\nabla_2f\\
    &=  \nabla_2V + \nabla_2R - R\nabla_2f\\
    &\phantom{=\ } -2\nabla_1\nabla_2f\nabla^1f + v^2{}_2\nabla_2f.
\end{align*}

\noindent Recall that on $\partial M$ we have that 
\begin{equation*}
    \nabla_1\nabla_\eta f = \hat{\nabla}_1 \nabla_\eta f + h_{11}\hat{\nabla}^1f,
\end{equation*}
where $\eta = -\partial_2$ is the inward pointing normal. Thus, replacing this in the equation above, we have that 
\begin{align*}
    \nabla_b v^b{}_2 + v^b{}_2\nabla_b f &=  \nabla_2V + \nabla_2R - R\nabla_2f\\
    &\phantom{=\ } + 2h(\hat{\nabla}f, \hat{\nabla}f) - 2\nabla_1(\partial_2 f)\nabla^1 f +  v^2{}_2\nabla_2f.
\end{align*}
From now on, we will set $\frac{\partial f}{\partial N} = H$ on $\partial M$. Taking this into account, using the equation above and that $\partial R/\partial N = HR - 2H'$ on $\partial M$, each boundary term in Equation (\ref{eq:variation}) vanishes except for the integrals of $2h(\hat{\nabla}f, \hat{\nabla}f)$ and $ 2\nabla_1(\partial_2 f)\nabla^1 f$. Notice that the integral of the latter also vanishes, because $\partial_2f = -H$ is (spatially) constant on $\partial M$. Since the entropy functional $\mathcal{W}_\infty$ is invariant under rescalings and diffeomorphisms, we have shown that 

\begin{theorem}[Entropy monotonicity formula]\label{th:2.4}
    Under the evolution equations
    \begin{equation}
        \begin{cases}
            \frac{\partial}{\partial t}g_{ab} = -2R_{ab} = -Rg_{ab},\\
            H(t) = \psi(t),\\
            \frac{\partial f}{\partial t} = -R - \Delta f + |\nabla f|^2 + \frac{1}{\tau},\\
            \frac{\partial f}{\partial N} = H, \hspace{10 pt} \text{on } \partial M\times [0,T)\\
            \frac{\partial \tau}{\partial t} = -1,
        \end{cases}    
    \end{equation}
    and the additional assumption that $H(t) \geq 0$, we have
    \begin{equation}\label{eq: monotonicity}
        \begin{split}
            \frac{d}{dt}\mathcal{W}_\infty &= \frac{1}{4\pi}\int_M2\left|R_{ab} + \nabla_a\nabla_bf - \frac{1}{2\tau}g_{ab}\right|e^{-f}d\mu\\
            &\phantom{=\ } + \frac{1}{4\pi}\int_{\partial M} 2H |\hat{\nabla} f|^2e^{-f}d\sigma \geq 0.
        \end{split}
    \end{equation}
\end{theorem}

\begin{remark}
    As a matter of fact, in both \cite{Lott} and \cite{Ecker} are shown entropy formulas for the functionals considered in them. Thus, the result above is a combination of their ideas applied to the case of the Ricci flow in compact surfaces with boundary.
\end{remark}


\section{Proof of Theorem 1.1}
\label{Sec:proof1.1}

In this section, we will give an outline of the proof for Theorem (\ref{th:1.1}). With this purpose in mind, let us recall a proposition 
due to \cite{Cortiss-Mur}. We must observe that, due to the hypotheses of Theorems (\ref{th:1.1}) and 
(\ref{th:1.2}), blow-up limits can be taken
as is shown in Sections 4 and 6 in \cite{Cortiss-Mur}.  We also refer the reader to \cite{Gianniotis}, where some compactness results are demonstrated in the case of manifolds with boundary.
\begin{proposition}\label{Prop: blow-up limits}
    Let $(M,g(t))$ be the solution to the Ricci flow (\ref{eq:Ricci flow}) and $[0,T)$ be the maximal interval of existence of this solution. Assume $H$ is bounded and there is an $\epsilon > 0$ such that on this interval $R > -\epsilon$. Hence, there are two possible blow-up limits for $(M,g(t))$ as $t\rightarrow T$. If the blow-up limit is compact, then it is a homothetically shrinking round hemisphere with totally geodesic boundary. While, if it is noncompact, it is (or its double is) a cigar soliton.
\end{proposition}

\begin{proof}
    Recall that the blow-up limit of a maximal solution $(M, g(t))$ to the Ricci flow (\ref{eq:Ricci flow}) is an ancient solution $(M_\infty, g_\infty(t), p_\infty)$ with totally geodesic boundary or no boundary at all; see \cite{Cortiss-Mur}. Then, in the first case, when the blow-up limit is compact, by doubling the manifold we obtain an ancient solution to the Ricci flow on a closed surface, and applying Theorem 26.1 of \cite{Hamilton} we conclude that the blow-up limit is a round hemisphere, since $R_\infty(p_\infty) > 0$. In the second case, the blow-up limit is noncompact; therefore, it is a singularity model of type II, and consequently it (or its double) is a cigar soliton, see Propositions 9.16 and 9.18 of \cite{Chow-Ni-Lu}.
\end{proof}

Now we state the corresponding version of Perelman's \textit{no local collapsing theorem} \cite{perelman2002} for the modified functional in the case of surfaces with boundary.

\begin{theorem}[No local collapsing theorem I]\label{th:3.3}
    Let $(M,g(t))$ be a maximal solution to the Ricci flow (\ref{eq:Ricci flow}) on a compact surface with boundary, with $t \in [0, T)$. Assume that the prescribed geodesic curvature $\psi$ is nonnegative. If $T < \infty$, then for any $\rho \in (0,\infty)$ there exists $k = k(g(0),T, \rho)>0$ such that the solution $g(t)$ is $k$-noncollapsed below the scale of $\rho$ as long as the solution exists.
\end{theorem}
\begin{proof}
    The idea of the proof is based on the arguments appearing in Theorem 5.35 in \cite{Chow-Ni-Lu}. Under the hypotheses above, we have that
    \begin{equation*}
        \mathcal{W}_\infty \geq \mathcal{W},
    \end{equation*}
    where $\mathcal{W}$ is the usual Perelman's entropy functional. Applying the logarithmic Sobolev inequality, we conclude that at $t = 0$ the functional $\mathcal{W}_\infty$ is bounded from below, \textit{i.e.},
    \begin{equation*}
        \mathcal{W}_\infty \geq \mathcal{W} \geq - C.
    \end{equation*}
    Since $M$ is compact, there is a $k_0$ such that $(M,g(t))$ is $k_0$-noncollapsed below the scale of $\rho$ for $t\leq T/2$, and since $T < \infty$,
    \begin{equation}\label{eq:8}
        \inf_{\tau \in [T/2, T + \rho^2]} \mu_\infty(g(0), \tau) := -C_1 > - \infty.
    \end{equation}
    When $t\in [T/2, T)$, then $t+ r^2\in [T/2, T + \rho^2]$ for any $r\leq\rho$ and using the monotonicity formula (\ref{eq: monotonicity}), we conclude that 
    \begin{equation}\label{eq:9}
        -C_1 \leq \mu_\infty(g(0), t + r^2) \leq \mu_\infty (g(t), r^2).
    \end{equation}
    In order to finish the proof, we will show that if $(M,g(t))$ were $k$-collapsed at the scale of $r$ at a time $t_0 \geq T/2$, then $\mu_\infty$ would be very large and negative. By definition, there is a point $p \in M$ such that for all $q\in B:= B(p,r)$ (at a time $t_0$), we have $|Rm (q)| \leq r^{-2}$ and $Vol(B)/r^2 \leq k$. Let $\Phi: = e^{-f/2}$, replacing this into the modified entropy functional we obtain the following.
    \begin{equation}\label{eq:10}
        \begin{split}
            \mathcal{W}_\infty &= \frac{1}{4\pi r^2}\int_M [(r^2R - 2\log \Phi - 2  )\Phi^2 + 4 r^2|\nabla \Phi|^2]d\mu\\
        &\phantom{=\ } + \frac{1}{4\pi}\int_{\partial M }2H \Phi^2d\sigma.
        \end{split}
    \end{equation}
    As suggested by Perelman in \cite{perelman2002}, let $\Phi(x) = e^{-c/2}\phi(d(p,x)/r)$, where $c$ is a normalization constant and $ \phi: \mathbb{R}^{\geq 0} \rightarrow \mathbb{R}$ is a smooth function equal to $1$ in $[0,1/2]$ decreasing monotonically to $0$ in $[1/2,1]$ and identically $0$ in $[1,\infty)$. Hence, it is well known that the first integral in Equation (\ref{eq:10}) is of the order of $\log k$. On the other hand, the second integral can be bounded as follows:
    \begin{equation*}
        \frac{1}{4\pi}\int_{\partial M}2H \Phi^2d\sigma < |\!| H |\!| Vol(\partial M \cap B).
    \end{equation*}
    Thus, the following inequality holds
    \begin{equation*}
          \mu_\infty(g(t_0), r^2) < \log k + |\!| H |\!| Vol(\partial M \cap B),
    \end{equation*}
    which is impossible by Equations (\ref{eq:8}) and (\ref{eq:9}).\\
\end{proof}

Recall that (for all $k>0$) the cigar is $k$-collapsed at some scale; therefore, we can now rule out the cigar soliton as the blow-up limit of singular solutions to the Ricci flow (\ref{eq:Ricci flow}) on a compact surfaces with boundary. To this end, we follow the arguments on Corollary 5.51 of \cite{Chow-Ni-Lu}. Suppose that $(M_\infty, g_\infty(t), p_\infty)$ is the blow-up limit of a solution $(M,g(t))$ to the Ricci flow (\ref{eq:Ricci flow}). Then, by Theorem (\ref{th:3.3}) we find that the initial manifold is $k$-noncollapsed at some scale $\rho$. Recall that the dilations $g_i$ of $g$ are defined by 
\begin{equation*}
    g_i(t) = \lambda_ig\left(t_i + \frac{t}{\lambda_i}\right),
\end{equation*}
where times $t_i$ and points $p_i$ are suitably chosen in such a way that $t_i \rightarrow T$ and 
\begin{equation*}
    \lambda_i = |R|_\text{max}(t) = R(p_i,t_i) \rightarrow \infty.
\end{equation*}
This implies that $(M_i,g_i(t))$ is $k$-noncollapsed below the scale of $\sqrt{\lambda_i}\rho$. Since $\lambda_i \rightarrow \infty$, we conclude that the blow-up limit $(M_\infty, g_\infty(t))$ is $k$-noncollapsed on all scales. In summary, we have shown the following.

\begin{corollary}[Precluding the cigar]
    Let $(M,g(t))$ be a compact surface with boundary which is a singular solution to the Ricci flow (\ref{eq:Ricci flow}) on a maximal interval $[0,T)$, with $R(t) \geq 0$ and $\psi\left(t\right) \geq 0$, then the cigar soliton cannot occur as its blow-up limit.
\end{corollary}

By Proposition (\ref{Prop: blow-up limits}) we conclude that the blow-up limit is therefore a shrinking round hemisphere with totally geodesic boundary. This implies that along the sequence of times $t_i\rightarrow T$ considered above we have that 
\begin{equation*}
    \lim_{i\rightarrow \infty}\frac{R_\text{max}(t_i)}{R_\text{min}(t_i)} = 1.
\end{equation*}
This gives a proof of Equation (\ref{eq:uniformized curvature}). Finally, following the arguments in Section 5 of \cite{Cortiss-Mur} completes the proof of Theorem (\ref{th:1.1}). Indeed, recall that the normalized Ricci flow is defined as follows. Given a solution $g(t)$ to the Ricci flow (\ref{eq:Ricci flow}) let $\Tilde{g}(\Tilde{t}):= \phi (t)g(t)$ where 
\begin{align*}
    \phi(t)A(t) = A(0) \hspace{5pt} \text{and} \hspace{5pt}
    \Tilde{t}(t) = \int_0^t \phi(\tau)d\tau.
\end{align*}
Therefore, the normalized Ricci flow exists for all time if $\int_0^T1/A(t) dt = \infty$. To show this, observe that 
\begin{equation*}
    A'(t) = -\int_MRd\mu = - 4\pi \mathcal{X}(M) + \int_{\partial M}2Hd\sigma \geq - 4\pi \mathcal{X}(M).
\end{equation*}
So, we obtain that $A(t) \leq 4\pi \mathcal{X}(M)(T - t)$, which proves the result. 
Now, the sequence of times $t_n \rightarrow T$ considered above, induces a sequence $\Tilde{t}_n \rightarrow \infty$. By the previous 
arguments, we know that the metrics $\lambda_ng(t_n)$ converge smoothly to a metric of constant curvature $\Bar{g}$ as $n\rightarrow \infty$. 
Define $c_n>0$ in such a way that $\Tilde{g}(\Tilde{t}_n) = c_n\lambda_ng(t_n)$; hence, $c_n\lambda_n = 1/A(t_n)$. We claim that the sequence 
$c_n$ is bounded. To see this, note that 
\begin{equation*}
    R_\text{max}(t)A(t) \geq \int_M Rd\mu = 4\pi\mathcal{X}(M) - \int_{\partial M} 2H d\sigma,
\end{equation*}
and since $H$ is bounded we have that 
\begin{equation*}
    -\int_{\partial M} H d\sigma \geq - |\!| H |\!| l(\partial M),
\end{equation*}
and since $R$ remains positive, the length of the boundary is decreasing in time. This implies there is a constant $K> 0$ such that 
\begin{equation*}
    R_\text{max}(t)A(t) \geq K.
\end{equation*}
Thus, $c_n = 1/(\lambda_nA(t_n))\leq 1/ K$, as we wanted to show. So, there is a subsequence $c_{n_k}$ converging to some $c$ and consequently the corresponding subsequence $\Tilde{g}(\Tilde{t}_{n_k}) = c_{n_k}\lambda_{n_k}g(t_{n_k})$ converges smoothly to the metric of constant curvature $c\Bar{g}$.


\section{Proof of Theorem 1.2}
\label{Sec:proof1.2}
Now we prove Theorem (\ref{th:1.2}). Thus, in this section
we let $(M, g(t))$ be a rotationally symmetric solution to the Ricci flow (\ref{eq:Ricci flow}) in the disk. From now on, we will only consider rotationally symmetric functions $f$ for the functional $\mathcal{W}_\infty$.

\begin{remark}
    From now on, $\mu_\infty(g,\tau)$ will represent the infimum of the functional with the additional restriction of $f$ being rotationally symmetric.
\end{remark}

\begin{theorem}[Entropy monotonicity formulla II]
    In the setup of the paragraph above, we have 
    \begin{equation}\label{eq:monotonicity II}
        \frac{d}{dt} \mathcal{W}_\infty = \frac{1}{4\pi}\int_M 2\left| R_{ab} + \nabla_a\nabla_bf - \frac{1}{2\tau}g_{ab}\right|e^{-f}d\mu \geq 0.
    \end{equation}
\end{theorem}
\begin{proof}
    First of all, it is important to mention that the rotational symmetry of $f$ is preserved under the equations 
    \begin{align*}
        \frac{\partial f}{\partial t} &= -\Delta f-R + \frac{1}{\tau},\\
        \frac{\partial f}{\partial N} &= H,
    \end{align*}
    corresponding to the evolution equation of $f$ under the flow  
    \begin{equation*}
        \frac{\partial}{\partial t}g_{ab} = -Rg_{ab} - 2\nabla_a\nabla_bf + \frac{1}{\tau}g_{ab},
    \end{equation*}
    which is equivalent to the Ricci flow. And secondly, observe that the term $$\frac{1}{4\pi}\int_{\partial M} 2H|\hat{\nabla}f|^2e^{-f}d\sigma$$ appearing in the monotonicity formula (\ref{eq: monotonicity}) vanishes since $f$ has rotational symmetry; therefore, $\hat{\nabla}f = 0$ in $\partial M$.
\end{proof}

Using arguments very similar to those employed in the previous section, 
we obtain a version for the \textit{no local collapsing theorem} 
in the case of rotational symmetry with no assumption on the sign of the prescribed geodesic curvature $\psi$. 

\begin{theorem}[No local collapsing theorem II] Let $g(t)$ be a rotationally symmetric maximal solution to the Ricci flow (\ref{eq:Ricci flow}) on the disk $M$ with $t \in [0,T)$. If $T< \infty$, then for any $\rho \in (0,\infty)$ there exists $k = k(g(0),T, \rho)>0$ such that the solution $g(t)$ is $k$-noncollapsed below the scale of $\rho$ as long as the solution exists.
\end{theorem}
\begin{proof}
     As before, we have that at $t = 0$ the functional $\mathcal{W}_\infty$ is bounded from below, there is a $k_0$ such that $(M, g(t))$ is $k_0$-noncollapsed below the scale of $\rho$ for $t \leq T/2$ and Equation (\ref{eq:8}) holds again \textit{i.e.}     \begin{equation}
        \inf_{\tau \in [T/2, T + \rho^2]} \mu_\infty(g(0), \tau) := -C_1 > - \infty.
    \end{equation}
    This time applying monotonicity formula (\ref{eq:monotonicity II}) we obtain that for $t\geq T/2$
    \begin{equation}\label{eq: 13}
        -C_1 \leq \mu_\infty(g(0), t + r^2) \leq \mu_\infty(g(t),r^2).
    \end{equation}
    As before, the proof finishes noticing that if $(M,g(t))$ were collapsing at some $t_0 \geq T/2$, then the functional would be very large and negative contradicting Equation (\ref{eq: 13}). Thus, let $x \in M$ such that $|Rm(p)| \leq r^{-2}$ for all $p\in B:= B(x,r)$ and 
    \begin{equation*}
        \frac{Vol(B)}{r^2} \leq k.
    \end{equation*}
    Observe that the function $\Phi$ used before is not necesarilly rotationally symmetric, unless $x$ is the origin of $M$. Hence, to fix this, instead of considering a function localized in the ball of radius $r$ around $x$ (the point where collapse is occuring), we will consider a function localized in the annulus $A$ of length at most $2r$ containing this ball. Indeed, let $O$ be the origin of $M$ and define $\Phi$ as follows. 
    \begin{equation*}
        \Phi(p) := e^{-c/2}\phi(|d(O,p) - d(O,x)|/r),
    \end{equation*}
    where $\phi$ is the same as before.
    Now we proceed to bound the modified entropy functional.
    \begin{equation*}
        \begin{split}
            \mathcal{W}_\infty &= \frac{1}{4\pi r^2}\int_M [(r^2R - 2\log \Phi - 2  )\Phi^2 + 4 r^2|\nabla \Phi|^2]d\mu\\
        &\phantom{=\ } + \frac{1}{4\pi}\int_{\partial M }2H \Phi^2d\sigma\\
        &= \frac{1}{4\pi r^2}\int_A[(r^2R - 2\log \Phi - 2)\Phi^2 + 4r^2|\nabla \Phi|^2]d\mu\\
        &\phantom{=\ } + \frac{1}{4\pi}\int_{\partial M \cap A}2H\Phi^2d\sigma.
        \end{split}
    \end{equation*}
    Recall that $r^2R$ is bounded in $B$ and due to the rotational symmetry it is also bounded in all of $A$. Observe that $|\nabla \Phi| = e^{-c/2}|\phi'|/r$; therefore, $4r^2|\nabla \Phi|^2 = 4e^{-c}|\phi'|$ is bounded. Now, $-2\log \Phi = c - \log \phi$, and recall that $\phi^2\log \phi \rightarrow 0$ as $\phi \rightarrow 0$ and is bounded far from $0$
    (this argument has been taken from \cite{Calegari}). 
    Finally, the boundary term is bounded exactly the same as before. Hence, we conclude that at $t = t_0$
    \begin{equation*}
        \mathcal{W}_\infty \leq c + constant.
    \end{equation*}
    To estimate the normalization constant $c$, we use that $\frac{1}{4\pi r^2}\int_M \Phi^2d\mu = 1$.
    Since $\phi \leq 1$ and it is localized in $A$ we have the following inequality
    \begin{equation*}
        1 \leq \frac{e^{-c}Vol(A)}{r^2}.
    \end{equation*}
    Note that there is a constant $C>0$ depending on $t_0$ such that
    \begin{equation*}
        Vol(A) \leq C Vol(B).
    \end{equation*}
    This implies that $c \leq \log (kC)$ and consequently, we have that 
    \begin{equation*}
        \mathcal{W}_\infty \leq \log (kC)+ constant,
    \end{equation*}
    which is not possible by Equation (\ref{eq: 13}).
\end{proof}

Recall that as mentioned in the previous section, under the hypotheses of Theorem (\ref{th:1.2}) we can take blow-up limits, regardless of the sign of the scalar curvature.
Therefore, applying the theorem above we can again rule out the cigar as a blow-up limit of $(M,g(t))$.

\begin{corollary}
    Let $(M, g(t))$ be the disk with a singular rotationally symmetric solution to the Ricci flow (\ref{eq:Ricci flow}). Then, the cigar soliton cannot be its blow-up limit.
\end{corollary}

Consequently, applying Proposition (\ref{Prop: blow-up limits}) once again, Equation (\ref{uniformized curvature 2}) holds under the hypotheses of Theorem (\ref{th:1.2}). To finish the proof, recall that in Section \ref{Sec:proof1.1} we have shown that when $\psi>0$, the normalized Ricci flow exists for all time; while, following the ideas in the proof of Theorem 1.2 in \cite{Cortiss-Mur} shows that the normalized Ricci flow exists for all time when $H < 0$. Thus, it remains to prove that the normalized Ricci flow converges to a metric of constant curvature and totally geodesic boundary up to a subsequence. This will follow if we are able to show that the sequence $c_n$ defined exactly the same as before is bounded. In the case in which $R>0$, this follows replicating the arguments of the previous section. While, if $R<0$ and $H>0$ we can consider a subsequence of times $t_i$ such that $\lambda_i$ is strictly increasing. Since $R$ is negative, the area of $M$ is increasing also, therefore 
\begin{equation*}
    |R|_\text{max}(t_n) A(t_n) \geq \lambda_0 A(t_0),
\end{equation*}
so that, $c_n= 1/(\lambda_n A(t_n))$ is bounded from above, as we wanted to show. Hence, we have shown subsequential convergence of the normalized Ricci flow to a metric of constant curvature, finishing the proof of Theorem (\ref{th:1.2}).

\bibliography{Referencias.bib}
    \bibliographystyle{alpha}

\end{document}